# MAGIC-VQA: Multimodal And Grounded Inference with Commonsense Knowledge for Visual Question Answering

### v1

@@ 6 Qualitative Analysis
Figure 5 compares MAGIC-VQA with GPT4o and Qwen2VL, demonstrating how our framework effectively integrates both explicit and implicit commonsense knowledge for enhanced visual question answering. As illustrated in Figure 5(a), while GPT4o struggles to deduce the complete answer (big buff ale) to input query, MAGIC-VQA successfully incorporate contextual knowledge, such as “Person X owns the tap sells beer” and “beverage dispenser used as a beer tap in a bar,” linking beer consumption, tap functionality, and beverage machines with the input question to arrive at the correct answer. In Figure 5(b), while Qwen2VL incorrectly identifies the colony as North Carolina, our MAGIC-VQA addresses this limitation by integrating explicit image-based commonsense knowledge about Virginia’s location, historical turnpikes, and wildlife, correctly concluding that the answer is Virginia. Confidence scores derived from implicit commonsense knowledge further reinforce the evidence for the final accurate prediction. Additional case studies are shown in Appendix B.
@@ Appendix A Additional Experiment Results
A.1 Ratio of Commonsense Knowledge Type
To determine the optimal ratio of commonsense knowledge types, we conduct an experiment to analyze how varying distributions of CS-PE, CS-EC, and CS-SI knowledge triplets impact the performance of Qwen2-VL on the ScienceQA dataset. Figure 6 indicates that the model achieves its highest accuracy of 75.11% with a 4:1:1 ratio, which places greater emphasis on CS-PE knowledge. Additionally, distributions favoring CS-PE triplets consistently result in improved performance. For example, a 3:1:2 ratio achieves an accuracy of 74.73%. In contrast, ratios prioritizing CS-EC or CS-SI, such as 1:2:3 or 1:3:2, yield lower accuracies of 72.56% and 72.41%, respectively. These results suggest that CS-PE is the most essential commonsense knowledge type for the ScienceQA, aligning with its focus on physical concepts and entities as discussed in Section 5.1. Notably, the 4:1:1 ratio closely mirrors the inherent distribution of commonsense knowledge in ScienceQA in Figure 3. This alignment suggests that tailoring the balance of commonsense knowledge to match the dataset’s inherent characteristics, as demonstrated by our approach, leads to the most significant performance improvements.
A.2 Number of Knowledge Triplets
@@
To optimize the integration of explicit commonsense knowledge into our model while minimizing the risk of introducing excessive noise, we investigate how varying the number of retrieved commonsense triplets $k$ impacts performance. We vary $k$ from 0 to 10, incrementally increasing the number of triplets provided to Qwen2VL andGPT4o-mini and measuring the corresponding accuracy. In Figure 7, the models achieve optimal performance when $k=6$ triplets are incorporated. Utilizing fewer than $k=6$ appears insufficient to provide the contextual information for efficient reasoning, while exceeding $k=6$ triplets includes irrelevant or redundant information, diminishing performance. Therefore, using $k=6$ represents the optimal balance/sweet spot for enriching the model with necessary knowledge while focusing on pertinent information.
A.3 Effect of Similarity Metric
@@
A.4 VLPM-Style Fine-Tuning Results
We further perform VLPM-style fine-tuning Han et al. (2020); Long et al. (2022b) to investigate the effectiveness of implicit multimodal commonsense in MAGIC-VQA framework. Table 4 presents a quantitative analysis of the impact of different input nodes—image (I), question (Q), and generated caption text (C)—on the performance of two baseline models (VILT and ALBEF) in VLPM-style fine-tuning. Each node type corresponds to a specific input modality, and when removed, the original embeddings from the pre-trained models are used instead of GCN-trained node embeddings. It is found that including all nodes consistently yields the best outcomes, underscoring the complementary contributions of visual, textual and caption inputs multimodal reasoning.
@@ Appendix D Baselines
  • LLaVA-1.6 (LLaVA-Next) Li et al. (2024): is an open-source Large Multimodal Model (LMM) designed for enhanced visual and conversational understanding built upon LLaVA Liu et al. (2024b). It supports higher input resolutions (up to 672x672 pixels) for finer visual detail recognition and incorporates improved visual instruction tuning for better reasoning and OCR capabilities. LLaVA 1.6 is highly efficient, using fewer than 1 million visual instruction tuning samples and a streamlined training process. Its versatility enables it to handle a wide range of applications, from image and text-based tasks to complex multimodal interactions, all while maintaining a minimalist and data-efficient design. We use llava-v1.6-mistral-7b-hf checkpoints for zero-shot testing.
@@ Appendix E Implementation Details
(a) ScienceQA
(b) TextVQA
(c) MMMU
Appendix F Retrieved Commonsense Triplet Cosine Similarity Distribution
Appendix G Concrete Example of Input Prompt
@@
Appendix H Commonsense Category Analysis Prompt Format

### v2

@@ 6 Qualitative Analysis
Figure 6 compares MAGIC-VQA with GPT4o and Qwen2VL, demonstrating how our framework effectively integrates both explicit and implicit commonsense knowledge for enhanced visual question answering. As illustrated in Figure 6(a), while GPT4o struggles to deduce the complete answer (big buff ale) to input query, MAGIC-VQA successfully incorporate contextual knowledge, such as “Person X owns the tap sells beer” and “beverage dispenser used as a beer tap in a bar,” linking beer consumption, tap functionality, and beverage machines with the input question to arrive at the correct answer. In Figure 6(b), while Qwen2VL incorrectly identifies the colony as North Carolina, our MAGIC-VQA addresses this limitation by integrating explicit image-based commonsense knowledge about Virginia’s location, historical turnpikes, and wildlife, correctly concluding that the answer is Virginia. Confidence scores derived from implicit commonsense knowledge further reinforce the evidence for the final accurate prediction. Additional case studies demonstrating the characteristics of each dataset and the benefit from each commonsense knowledge injection component are shown in Appendix B.
@@ 7 Conclusion
Acknowledgment
This study was supported by funding from the Google Award for Inclusion Research Program (G222897).
@@ Appendix A Additional Experiment Results
A.1 Number of Knowledge Triplets
@@
Inspired by Wu et al. (2024), to optimize the integration of explicit commonsense knowledge into our model while minimizing the risk of introducing excessive noise, we investigate how varying the number of retrieved commonsense triplets $k$ impacts performance. We vary $k$ from 0 to 10, incrementally increasing the number of triplets provided to Qwen2VL andGPT4o-mini and measuring the corresponding accuracy. In Figure 7, the models achieve optimal performance when $k=6$ triplets are incorporated. Utilizing fewer than $k=6$ appears insufficient to provide the contextual information for efficient reasoning, while exceeding $k=6$ triplets includes irrelevant or redundant information, diminishing performance. Therefore, using $k=6$ represents the optimal balance/sweet spot for enriching the model with necessary knowledge while focusing on pertinent information.
A.2 Effect of Similarity Metric
@@
A.3 VLPM-Style Fine-Tuning Results
We further perform VLPM-style fine-tuning Long et al. (2022b); Li et al. (2025b); Wang et al. (2025c) to investigate the effectiveness of implicit multimodal commonsense in MAGIC-VQA framework. Table 6 presents a quantitative analysis of the impact of different input nodes—image (I), question (Q), and generated caption text (C)—on the performance of two baseline models (VILT and ALBEF) in VLPM-style fine-tuning. Each node type corresponds to a specific input modality, and when removed, the original embeddings from the pre-trained models are used instead of GCN-trained node embeddings. It is found that including all nodes consistently yields the best outcomes, underscoring the complementary contributions of visual, textual and caption inputs multimodal reasoning.
@@ A.3 VLPM-Style Fine-Tuning Results
A.4 Additional Results on Knowledge Intensive Benchmarks
We further attach the result of MAGIC-VQA on two knowledge-intensive tasks including VCR and A-OKVQA as a further complementary analysis.
Figure 9: Further Experiment results on VCR and A-OKVQA
@@ Appendix D Baselines
  • LLaVA-1.6 (LLaVA-Next) Li et al. (2024a): is an open-source Large Multimodal Model (LMM) designed for enhanced visual and conversational understanding built upon LLaVA Liu et al. (2024b). It supports higher input resolutions (up to 672x672 pixels) for finer visual detail recognition and incorporates improved visual instruction tuning for better reasoning and OCR capabilities. LLaVA 1.6 is highly efficient, using fewer than 1 million visual instruction tuning samples and a streamlined training process. Its versatility enables it to handle a wide range of applications, from image and text-based tasks to complex multimodal interactions, all while maintaining a minimalist and data-efficient design. We use llava-v1.6-mistral-7b-hf checkpoints for zero-shot testing.
@@ Appendix E Implementation Details
Appendix F Concrete Example of Input Prompt
@@
Appendix G Commonsense Category Analysis Prompt Format
@@
Appendix H Retrieved Commonsense Triplet Cosine Similarity Distribution
(a) ScienceQA
(b) TextVQA
(c) MMMU